OpenStudy (anonymous):

find the derivative of the function h(x)=sqrt(2-x)

myininaya (myininaya):

use chain rule

myininaya (myininaya):

by the sqrt( ) can be written as ( )^(1/2)

OpenStudy (anonymous):

we are just in the first section and haven't learned the chain rule, and are required to use a longer method

myininaya (myininaya):

ok the definition then

OpenStudy (anonymous):

right

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h} =f'(x)\]

myininaya (myininaya):

where we have f(x)=sqrt(2-x) and so f(x+h)=sqrt(2-(x+h)) <--here just replaced my x's with (x+h)'s

myininaya (myininaya):

now plug in

myininaya (myininaya):

sometimes a good method when you are trying to evaluate a limit algebraically and you have a difference/sum of radicals rationalizing might help

myininaya (myininaya):

to rationalize recall conjugates

myininaya (myininaya):

for example the conjugate of a-b is a+b and the conjugate of a+b is a-b

myininaya (myininaya):

do you want to try to see what happens if you do the following: \[\lim_{h \rightarrow 0} \frac{\sqrt{2-(x+h)}-\sqrt{2-x}}{h} \cdot \frac{\sqrt{2-(x+h)}+\sqrt{2-x}}{\sqrt{2-(x+h)}+\sqrt{2-x}}\]

OpenStudy (anonymous):

right, I got it down to -1/sqrt2-x-h +sqrt2-x

myininaya (myininaya):

one sec let me check

myininaya (myininaya):

that looks pretty sweet

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{-1}{\sqrt{2-x-h}+\sqrt{2-x}}\]

myininaya (myininaya):

your job is to just replace h with 0

myininaya (myininaya):

any you could do a little magic afterwards to make the answer look a bit more tidy

OpenStudy (anonymous):

so -1/2sqrt2-x?

OpenStudy (anonymous):

thank you!

myininaya (myininaya):

\[\frac{-1}{2 \sqrt{2-x}} \text{ is \right!} \]